The critical value or t-value of a one-tailed test when n=18 for 95% of confidence interval is 1.734 *


1
Expert's answer
2021-06-10T04:56:41-0400
"df=n-1=18-1=17, \alpha=0.05"

The critical value or t-value of a one-tailed test when n=18 for 95% of confidence interval is "1.739607\approx1.734" for right-tailed test.


The critical value or t-value of a one-tailed test when n=18 for 95% of confidence interval is 1.734 (right-tailed).